Calculate Ka for 1 M NaCH3COO

pH = 9.37
H+ = 4.203 X 10^-10
OH- = 2.402 X 10^-5
CH3COO- = 1.00 X 10^0
CH3COOH = 2.402 X 10^-5

________________
Calculate Ka using the formula
Kw = Ka X Kb and the appropriate equilibrium constant for 1 M NaCH3COO.
NaCH3COO ---> Na+ (neutral) + CH3COO-.
Is the conjugate base for CH3COOH.

CH3COOH Ka1 = 1.8 X 10^-5
______________________

Calculate Ka for 1 M NH4Cl

pH = 4.26
H+ = 2.336 X 10^-5
OH- = 4.267 X 10^-10
NH4+ = 1.00 X 10^0
NH3 = 2.366 X 10^-5

_____________________
Calculate Ka using the formula
Kw = Ka X Kb and the appropriate equilibrium constant for NH4Cl
NH4Cl = conjugate acid NH3..

NH3 Kb1 = 1.76 X 10^-5



Please show me how you got these ... Thank you
Answer #1

Ka, CH3COOH = [H+][ CH3COO- ]/[ CH3COOH ] = 1.75x 10-5

………………….

You have to calculate Kb of CH3COO-, not a Ka.

Kb = 10-14 / Ka,CH3COOH = 10-14 / 1.8 X 10-5 = 5.5x10-10

……………………..

Ka,NH4+ = [NH3][ H+]/[ NH4+ ] =

            = 2.366 X 10^-5 x 2.366 X 10^-5 / 1.00 X 10^0 =

             = 5.60 x 10-10

………………………….

You have to calculate Ka,NH4+ using the value of Kb,NH3

Ka,NH4+ = 1x10-14 / Kb,NH3 = 1x10-14 / 1.76 X 10^-5 = 5.68 x10-10

Note: Kw = 1x10-14